OpenStudy (firejay5):

Heather and Matthew take 2,040 seconds to walk eastward along a straight road to a store 2.0km away. What's is the average velocity? *Use the following below in your answer for identification issues* V = D = T = *Show and explain work please!!!!!!!!*

OpenStudy (firejay5):

@clamin

OpenStudy (anonymous):

Average velocity, in words, is just the representation of the distance traveled and how long it took to travel said distance. In math, this is depicted by division. By taking the total distance traveled and dividing it by the time taken, you get a speed. distance is measured in some form of meters (m) time is measured in seconds (s) velocity is measured in (distance / time) = meters per seconds (m/s) In this case, you know distance D and you know time T. So it is as simple as dividing the two numbers, keeping track of units. V = D / T = (2 km) / (2040 s) = 0.00098 km/s If we multiply by the fraction (1000 m / 1 km), we can get a more easily understandable and usable answer: V = (0.00098 km/s) * (1000 m / 1 km) = 0.98 m/s

OpenStudy (firejay5):

how did you get 0.00098km/s @Pompeii00

OpenStudy (anonymous):

Take 2 km and divide it by 2,040 s 2 / 2040 = 9.8 x 10^-4 = 0.00098 km/s